OpenStudy (anonymous):

I need to understand how to solve this:

OpenStudy (anonymous):

\[(3x+6) \div (x ^{2}+3x+2)) \times ((x-1) \div (2x+8))\]

OpenStudy (anonymous):

I need to simplify this

OpenStudy (anonymous):

this looks easy, but then I just dont seem to get the right answer... strange

OpenStudy (anonymous):

Whenever you see something like this, you have to (1) try and factor the polynomials (2) take out any common factors in factors Here,\[\frac{3x+6}{x^2+3x+2} \times \frac{x-1}{2x+8}=\frac{3(x+2)}{(x+2)(x+1)}\times \frac{x-1}{2(x+4)}\]In the first part, you have a common factor of (x+2) in the numerator and denominator. These will divide to 1 (i.e. they cancel), so you're left with\[\frac{3}{x+1}\times \frac{x-1}{2(x+4)}=\frac{3(x-1)}{2(x+1)(x+4)}\]

OpenStudy (anonymous):

I don't think it can be simplifed

OpenStudy (anonymous):

I got stuck too ^^"

OpenStudy (anonymous):

I just hate these problems.... integration and derivation is even simpler to me than these... :( I guess my brain developed in such a way that I am dumb maths-wise...

OpenStudy (anonymous):

when you say divide do you mean this:\[(3x+6)/(x^2+3x+2)\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lol, it happens, "alot"

OpenStudy (anonymous):

calm down... ;p

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

(3x+6)=3(x+2) and x^2+3x+2=(x+2)(x+1). Dividing the two you get, 3/(x+1)

OpenStudy (anonymous):

I am calm, guys, it is just that I never took time to learn maths proparly, during the last 5 years, I haven't opened a maths textbook at home, and used to day dream in maths classes. This month is the first time in my life that I actually started to study maths at home

OpenStudy (anonymous):

but why are you dividing the 2? .-. aren't we supposed to multiply?

OpenStudy (anonymous):

and that's good, you should be proud.

OpenStudy (anonymous):

loki, why did you divide both the fractions when you're supposed to multiply?

OpenStudy (anonymous):

it is good, but I am a few years late.... I used to spend ALL my time learning Computer Science stuff.... like programming (I know fluently C++, C, Assembly, Pascal, and some others), metaphysics...

OpenStudy (anonymous):

damn, I should have listened to my mom

OpenStudy (anonymous):

nah, you're no late, if you were too, then you won't understand this math you're seeing now :)

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

I don't think I did divide both fractions.

OpenStudy (anonymous):

listened to your mom about what?

OpenStudy (anonymous):

don't get caught up in the 'few years late' thing

OpenStudy (anonymous):

those are the things that get into your head when you are stressed.

OpenStudy (anonymous):

go sky diving , you'll need it

OpenStudy (anonymous):

LOL, back in the days, at school, I used to be quite careless about school.... you know.. never studying at home and stuff... and i remember how my mother used to tell me that I will be sorry :) and I am now :) but no, mimi, i am not stressed at all... I am calm

OpenStudy (anonymous):

stress has nothing to do with calmness

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

you just have too much in your head right now and you need to filter them all out

OpenStudy (anonymous):

thinking too much, just too much, go jog or do some sort of activity to stop that thinking

OpenStudy (anonymous):

rest that head of yours

OpenStudy (anonymous):

alright... maybe I will stop maths for a while... to clear the mind, and go do some physics instead (i love physics)

OpenStudy (anonymous):

NO! lol

OpenStudy (anonymous):

no more studying for now,

OpenStudy (anonymous):

I think you should keep studying

OpenStudy (anonymous):

alright, I got a virtual mom and dad, now, whom should I listen to ... LOL

OpenStudy (anonymous):

loki >_> he's been studying for how long now? 2 weeks? time for him to chill a bit

OpenStudy (anonymous):

not good, he must have fun the rest of the day :)

OpenStudy (anonymous):

the exam's on tuesday

OpenStudy (anonymous):

he is chilled

OpenStudy (anonymous):

so? what's the point of him studying with that mood of his?

OpenStudy (anonymous):

nope he isn't, when I say chill, I say go out and have FUN

OpenStudy (anonymous):

is your mood alright, oktal?

OpenStudy (anonymous):

he must rest

OpenStudy (anonymous):

I am smiling right now, and I think I dont feel like going to do anything else but study, I feel good enough to keep pumping info to my head, so I thing i can just swich to physics and then after some time back to maths... :)

OpenStudy (anonymous):

think*

OpenStudy (anonymous):

>_>... you're still going to have the day off before the big day

OpenStudy (anonymous):

and then chemistry and then English and repeat...

OpenStudy (anonymous):

._. loki you're destroying him, it's too much for him

OpenStudy (anonymous):

i'm teasing ;p

OpenStudy (anonymous):

>_>...

OpenStudy (anonymous):

yeah, starlet, I will take a monday off... and I actually dont have to study English as it will be an essay... so nothig special, I will just have to concentrate so I dont do stupid spelling mistakes :)

OpenStudy (anonymous):

Good, reached an agreement ^_^

OpenStudy (anonymous):

Yes, and make sure you have a good essay structure too.

OpenStudy (anonymous):

you promised me, don't forget that andy

OpenStudy (anonymous):

ok, mom

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

I'm too young to be your mum

OpenStudy (anonymous):

is that a furious face?

OpenStudy (anonymous):

I don't give furious faces, not my style

OpenStudy (anonymous):

LOL just remembered my angry face LOL

OpenStudy (anonymous):

don't even think abt it

OpenStudy (anonymous):

wait, what was it? something like..

OpenStudy (anonymous):

don't >_>

OpenStudy (anonymous):

something, hmm.. oh I remember...

OpenStudy (anonymous):

oktal, i like how you've had that account for three seconds and you already have 11 fans

OpenStudy (anonymous):

ANDRIUS!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

loki, what does this face mean? :@

OpenStudy (anonymous):

>_>.............

OpenStudy (anonymous):

i have no idea. it looks like a reindeer to me

OpenStudy (anonymous):

LOL =D

OpenStudy (anonymous):

ask google and he'll tell you

OpenStudy (anonymous):

anyway, it means angry face

OpenStudy (anonymous):

that?

OpenStudy (anonymous):

and, starlet, Google is SHE

OpenStudy (anonymous):

it looks soft

OpenStudy (anonymous):

right, thought it was the happy relaxed sorta face ~

OpenStudy (anonymous):

Google = HE

OpenStudy (anonymous):

well, it has another meaning too, FRENCH KISS

OpenStudy (anonymous):

rofl, you come up with the most random stuff

OpenStudy (anonymous):

right, that wasn't your conclusion angoo

OpenStudy (anonymous):

he got it from google >_>

OpenStudy (anonymous):

NO, I am serious, it really means that

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

I know you're serious, that's what's funny

OpenStudy (anonymous):

so you are loughing that I am certain that :@ means french kiss?

OpenStudy (anonymous):

it's not

OpenStudy (anonymous):

laughing :D

OpenStudy (anonymous):

if so, then you weren't showing any happy relaxed faces last time andrius >_>

OpenStudy (anonymous):

I don't find it funny, both of you

OpenStudy (anonymous):

-.- relaxed face

OpenStudy (anonymous):

LOL =D

OpenStudy (anonymous):

that's not a relaxed face

OpenStudy (anonymous):

Arman does that face all the time

OpenStudy (anonymous):

anyway, it is a good thing to give someone a kiss, isn't that?